# Growth, discovery and characterization of single crystalline Eu0.8Pt6Al15.8

### v1

@@ Abstract
We report the discovery of a ternary compound, Eu0.8Pt6Al15.8. We determine its chemical and structural characteristics based on energy-dispersive X-ray spectroscopy as well as both powder and single-crystal X-ray diffraction, demonstrating that it crystallizes in a hexagonal structure type EuPt6Al16 with no reported structural analog. The electronic and magnetic properties are characterized by temperature- and field-dependent magnetization, and temperature-dependent resistance measurements, revealing that the Eu2+ magnetic moments order antiferromagnetically below 2.8 K.
@@
organization=Institut für Anorganische und Analytische Chemie, Universität Münster,addressline=Corrensstrasse 20, city=Münster, postcode=48149, country=Germany
@@ 1 Introduction
Recently, a Eu2Pt6Al15 phase was reported to exhibit a valence transition from Eu2+ to Eu3+ upon cooling below 45 K [11], according to Mössbauer spectroscopy, resistance and magnetic susceptibility measurements. This phase was obtained in polycrystalline form by arc-melting Eu, Pt and Al together, followed by a thermal treatment. As part of an attempt of obtaining that phase in single crystalline form by solution growth [12], we discovered a different ternary: Eu1-xPt6Al16-y, with $x\sim 0.2$ and $y\sim 0.2$. Since it adopts a structure that has not been reported in any other compound, in this paper we report the details for obtaining single crystals of this phase, as well as its basic structural, chemical, and magnetic characterization.
@@ 2 Experimental details
Single crystals of Eu0.8Pt6Al15.8, were obtained by the high-temperature solution growth method [13] out of a ternary melt with an initial nominal composition of Eu2Pt10Al88. The starting materials were elemental Pt powder (Ames National Laboratory, 99.9+% purity), pieces of Al rod (Alfa Aesar, 99.999% purity) and Eu pieces (Materials Preparation Center - Ames National Laboratory, 99.99+% purity).
There is a large exothermic reaction between Al and Pt/Pd upon melting from separate elements into a single-phase liquid. To prevent this exothermic reaction from occurring inside the alumina crucible in which the crystals grow, and potentially cracking the former, a step similar to that reported for the growth of the Al-Pd-Mn single grain quasicrystal [14] was followed. This consisted of pre-melting the Al and Pt using an arc furnace. Pt powder was first pressed into a pellet, arc-melted, weighed and then arc-melted in contact with Al pieces, to form a single button with composition of Pt10Al88. The arc-melted button was placed in a 2 ml alumina fritted Canfield Crucible Set [15, 16] together with the corresponding amount of Eu to achieve the desired initial composition. The crucible with the elements was sealed in a fused silica tube with a partial pressure of $\sim 1/6$ atm of Ar and placed inside a box furnace. It was heated to 1180 C, kept at that temperature for 5 hours, and slowly cooled down to 900 C over the course of 96 hours. At 900 C the excess solution was decanted with the aid of a centrifuge [13]. Pictures of the relatively large, obtained crystals are shown in Fig. 1(b).
Figure 1: Crystal structure of Eu0.8Pt6Al15.8, with the Eu atoms in blue, the Pt atoms in dark gray and the Al atoms in light gray: (a) including polyhedra around the Eu atoms in the $2c$ sites without vacancies (blue) and around the Eu atoms in the $2e$ sites with vacancies (cyan). (b) Photograph of one of the obtained single crystals. (c) Unit cell viewed down the $c$-axis; (d) unit cell viewed down the $a$-axis.
The concentration levels of each element were determined by energy dispersive X-ray spectroscopy (EDS) quantitative chemical analysis using an EDS detector (Thermo NORAN Microanalysis System, model C10001) attached to a JEOL scanning-electron microscope (SEM). An acceleration voltage of 16kV16\ $\text${kV}, working distance of 10mm10\ $\text${mm} and take-off angle of $35^{\circ}$ were used to measure all standards and crystals with unknown composition. Since we were able to obtain polycrystalline arc-melted and single-crystalline samples with the reported composition of Eu2Pt6Al15 [11, 12], flat samples of this phase were used as a standard for Eu, Pt and Al quantification. The spectra were fitted using NIST-DTSA II Microscopium software [17]. The composition of each crystal was measured at four different positions on an arbitrary crystal faces. The average compositions and error bars were obtained from these data, accounting for both inhomogeneity and goodness of fit of each spectra.
@@
For single-crystal X-ray diffraction (SCXRD) measurements, regularly shaped crystal fragments of Eu0.8Pt6Al15.8 were obtained by fracturing the flux grown crystals. The fragments were glued to quartz fibers using beeswax. The data sets were collected on a Stoe IPDS-II diffractometer (graphite monochromatized Mo $K\alpha$ radiation; oscilation mode). Numerical absorption correction [19] was applied to the data series.
@@
The temperature-dependent, AC resistance of the samples was measured using a Quantum Design Physical Property Measurement System (PPMS) using the AC transport (ACT) option, with a frequency of 17Hz17\ $\text${Hz} and a 3mA3\ $\text${mA} excitation current. The resistance both parallel and perpendicular to the $c$-axis was measured using a standard four-contact geometry. Electrical contacts with less than $1.5\ \Omega$ resistance were achieved by spot welding $25\ \mu\text{m}$ Pt wire to the samples, followed by adding Epotek H20E silver epoxy, and curing the latter for 1 hour at $120^{\circ}\text{C}$.
@@ 3.1 Chemical and structural characterization
Table 1: Single crystal X-ray diffraction measurement and refinement information for Eu0.8Pt6Al15.8.
Chemical formula Eu0.795(7)Pt6Al15.81(6)
Formula Weight 1717.6
@@
Data/restraint/parameters 2204/0/48
Goodness-of-fit on $F^{2}$ 1.8578
Final $R$ indexes $I\geq 2\sigma(I)$] $R_{1}=0.0304$, $wR_{2}=0.0674$
Final $R$ indexes [all data] $R_{1}=0.0483$, $wR_{2}=0.0675$
Largest diff. peak/hole 14.09 $e^{-}\text{\AA }^{-3}$/-6.15 $e^{-}\text{\AA }^{-3}$
Single-crystal X-ray diffraction measurement on the obtained crystals of Eu0.8Pt6Al15.8, revealed that it adopts a crystal structure Eu1-xPt6Al16-y ($x\sim 0.2$, $y\sim 0.2$) that does not match with any currently reported one. Table 1 shows the information on the measurement and refinement. The system is hexagonal and belongs to the space group $P\bar{6}2m$. A diagram of the unit cell basis of the structure is presented in Fig. 1, and the detailed information of the atomic positions, occupancies and thermal displacement parameters are given in Tab. 2. There are two distinct Eu sites ($2c$ and $2e$) located in the center of the blue and cyan polyhedra defined by the surrounding Pt and Al atoms in Fig. 1(a). The Eu $2e$ sites (in the cyan polyhedra) have an occupancy of around 59(1)%, whereas the other has full occupancy. Besides the Eu1 position, also the Al10 position shows an occupancy lower than 1, while all other positions are fully occupied within three standard deviations. The resulting composition is therefore Eu0.795(7)Pt6Al15.81(6), corresponding to a $x=0.205(7)$ and a $y=0.19(6)$. Since the Eu occupation is close to 1/2, different symmetry reductions were tested, however, no conclusive structural model was obtained in a lower symmetric space group. A potential symmetry loss, however, might be indicative by the remaining residual electron density.
Table 2: Refined atomic positions, site occupancy, and isotropic thermal displacement parameters for Eu0.8Pt6Al15.8, according to SCXRD.
@@
Some of the crystals obtained were ground in order to measure powder X-ray diffraction, obtaining the pattern shown in Fig. 2, plotted with black symbols. A Rietveld refinement was performed to fit this pattern, yielding $\chi^{2}=4.5$ and $R_{w}=15.5$. The model structure obtained from the single crystal X-ray refinement was used as the starting point; only the global factor, lattice parameters, atomic positions, microstrains, sample displacement and background were refined. The rest of the parameters, including the thermal displacements and the atomic fractions were kept fixed according to the values reported for the single crystal. The fitted curve is plotted as a red line, showing good agreement with the measured data, as also indicated by the difference (green line).
Figure 2: Powder X-ray diffraction pattern measured for Eu0.8Pt6Al15.8 (black symbols), the best fit obtained by Rietveld refinement (red line) and the residues (green line).
Table 3: Elemental analysis of Eu0.8Pt6Al15.8.
@@ 3.2 Electronic and magnetic characterization
Fig. 3 shows the ZFC and FC temperature-dependent magnetization normalized by the applied field of 50 Oe, applied along three mutually orthogonal directions: along the $a$-axis (in blue), along the $c$-axis (in green), and along the reciprocal lattice vector $b^{*}$ (in green). These three directions are schematically indicated with arrows on a prism drawing in the main panel. The inset shows an enlarged view of the low temperature range of these results, in order to see the feature more clearly in all three directions of applied field. The magnetization along $a$ and $b^{*}$ shows a downturn upon cooling below 3 K, whereas the magnetization along $c$ shows a more subtle kink and remains higher at low temperatures. The ZFC and FC data remain indistinguishable from each other throughout the measured temperature range.
Figure 3: Main panel: temperature-dependent magnetization of Eu0.8Pt6Al15.8 normalized by the applied field of 50 Oe, oriented along $a$ (blue symbols), $b^{*}$ (red symbols) and $c$ (green symbols). All the shown results were done following both ZFC and FC protocls, which cannot be appreciated in the plot because they coincide with each other. A schematic diagram indicating the different directions with respect to the unit cell was also included. Inset: enlarged view of the low temperature range in order to visualize the feature in greater detail.
@@
The magnetization, in units of Bohr magneton, $\mu_{B}$, per Eu atom is plotted as a function of the applied field along $a$, $b^{*}$ and $c$ in Fig. 4, measured at a constant temperature of 2 K with an increasing magnetic field. As was the case for Fig. 3, the formula unit of Eu0.795(7)Pt6Al15.81(6) was used for computing the intrinsic value of magnetization per Eu. The inset shows an enlarged view of the low-field data, showing a linear behavior that extrapolates to $M(H=0)=0$. No hysteresis is expected in any of the three directions, since no difference was observed between the temperature-dependent ZFC and FC measurements done at 50 Oe (see Fig. 3). As a reference, the black squares in the inset of Fig. 4 show the corresponding ZFC and FC magnetization values extracted from the $M(T)$ measurements at 2 K and a field of 50 Oe applied parallel to $c$, both consistent with the $M(H)$ results. The measurements along $a$ and $b^{*}$ both display some metamagnetic transition around 12 kOe, which is broadened due to the fact that the measurements were taken at a temperature that is close to the transition temperature. Combined, the $M(T)$ and $M(H)$ measurements shown in Figs. 3 and 4 indicate that the sample at low temperatures exhibits antiferromagnetic ordering, without signs of any ferromagnetic component.
@@
The magnetization shown in Fig. 4 does not reach full saturation at the maximum field measured, but it is consistent with a saturation of 7 $\mu_{B}$ expected for Eu2+ ions with $J=7/2$, indicated with a horizontal dashed line in the figure when using the proper stoichiometry of Eu0.795(7)Pt6Al15.81(6). The $\sim 5\%$ difference between the magnetization for $H||a$ (blue curve) and $H||b^{*}$ (red curve) could be due to the background subtraction varying slightly when the Kel-F disk axis is oriented vertically or horizontally  [^1] The background signal was measured independently only for the disk oriented with its axis horizontally (perpendicular to the axis of the straw).
@@
Figure 4: Main panel: Magnetization of Eu0.8Pt6Al15.8 as a function of the applied field, oriented along $a$ (blue), $b^{*}$ (red) and $c$ (green), measured at 2 K. The horizontal dashed line indicates the expected saturation magnetization for Eu2+ ions. Inset: Enlarged view of the low-field results. The vertical dashed line indicates the field at which the $M(T)$ measurements were performed, and the overlapping black squares show the corresponding ZFC and FC magnetization at 2 K obtained from the $M(T)$ measurements done with a field of 50 Oe applied parallel to $c$.
@@
Figure 5: Low temperature polycrystalline average of the magnetic susceptibility of Eu0.8Pt6Al15.8 $\chi_{\text{ave}}$ (black, left axis), and $d(\chi_{\text{ave}}T)/dT$ (magenta, right axis) as a function of temperature.
Figure 6: Main panel: Normalized resistance of Eu0.8Pt6Al15.8 as a function of temperature, measured upon cooling and warming, for the applied current oriented along $a$ (blue) and $c$ (green). Insets: enlarged view of the low temperature range for both measurements; the black arrows indicate the estimated transition temperature according to $d(\chi_{\text{ave}}T)/dT$.
Figure 7: Temperature-dependent magnetization of Eu0.8Pt6Al15.8 normalized by the applied field of 100 Oe applied along the $c$-axis (solid black symbols, left axis), and its inverse (open magenta symbols, right axis). The blue line represent the linear fit applied according to the Curie-Weiss law, in order to obtain the effective magnetic moment, $\mu_{\text{eff}}$, and the Weiss temperature, $\theta_{CW}$.
@@
This yielded $\mu_{\text{eff}}=8.0(1)\ \mu_{B}$, which agrees with the expected value for Eu2+ ions (7.94 $\mu_{B}$) within uncertainty; and $\theta_{CW}=-5.0(1)$ K, which would indicate the paramagnetic phase has predominantly antiferromagnetic interactions, and is consistent with the low-temperature antiferromagnetic ordering.

### v2

# Growth, discovery and characterization of single crystalline Eu0.8Pt6Al16.4

@@ Abstract
We report the discovery of a ternary compound, Eu0.8Pt6Al16.4. We determine its chemical and structural characteristics based on energy-dispersive X-ray spectroscopy as well as both powder and single-crystal X-ray diffraction, demonstrating that it crystallizes in a hexagonal structure type EuPt6Al17 with no reported structural analog. The temperature- and field-dependent magnetization, and temperature-dependent resistance measurements, reveal that the Eu2+ magnetic moments order antiferromagnetically below 2.8 K.
@@
organization=Institut für Anorganische und Analytische Chemie, Universität Münster,addressline=Corrensstrasse 28, city=Münster, postcode=48149, country=Germany
@@ 1 Introduction
Recently, a Eu2Pt6Al15 phase was reported to exhibit a valence transition from Eu2+ to Eu3+ upon cooling below 45 K [11], according to Mössbauer spectroscopy, resistance and magnetic susceptibility measurements. This phase was obtained in polycrystalline form by arc-melting Eu, Pt and Al together, followed by a thermal treatment. As part of an attempt of obtaining that phase in single crystalline form by solution growth [12], we discovered a different ternary: Eu1-xPt6Al17-y, with $x\sim 0.2$ and $y\sim 0.6$. Since it adopts a structure that has not been reported in any other compound, in this paper we report the details for obtaining single crystals of this phase, as well as its basic structural, chemical, and magnetic characterization.
@@ 2 Experimental details
The crystals of Eu0.8Pt6Al16.4 were grown by a similar method as described in Ref. 12, with a different initial ratio of Eu:Pt:Al of 2:10:88. Platinum powder with purity of 99.9+%, aluminum rod with purity of 99.999% and europium pieces with purity of 99.99+%. The platinum and aluminum were melted together in an arc-furnace [12] in the ratio 10:88. We placed the resulting button and the missing amount of europium in a 2 ml Canfield Crucible Set [13]. We used a box furnace to heat the crucible with the elements sealed inside a silica ampule filled with 0.2 atm of argon. It was first kept for 5 hours at 1180 C, and then cooled down to 900 C over 96 hours. At this temperature, we decanted the solution using a centrifuge [14], obtaining relatively large crystals as the one shown in Fig. 1(b).
Figure 1: Crystal structure of Eu0.8Pt6Al16.4, with the Eu atoms in blue/cyan, the Pt atoms in dark gray and the Al atoms in pink: (a) including polyhedra around the Eu atoms in the $2c$ sites without vacancies (blue) and around the Eu atoms in the $2e$ sites with vacancies (cyan). (b) Photograph of one of the obtained single crystals. (c) Unit cell viewed down the $c$-axis; (d) unit cell viewed down the $a$-axis.
The elemental composition of the crystals was determined by energy dispersive X-ray spectroscopy (EDS) using a JOEL scanning-electron microscope with a Thermo NORAN Microanalysis detector. Flat samples of Eu2Pt6Al15 [11, 12] were used as a standard for Eu, Pt and Al quantification. The composition of each crystal was measured at four different positions on arbitrary crystal faces. All the measurements were done with a 16kV16\ $\text${kV} acceleration voltage, 10 mm working distance, and $35^{\circ}$ take-off angle. We used the NIST-DTSA II Microscopium software [15] to fit each spectrum, and eventually obtain average compositions and error bars account for the goodness of fit as well as inhomogeneity in the composition.
@@
Magnetization measurements were performed with a Quantum Design Magnetic Property Measurement System (MPMS Classic and MPMS3) device based on a superconducting quantum interference magnetometers, covering 1.8 K $\leq T\leq$ 300 K and fields up to $|H|=70$ kOe. Each crystal was oriented by Laue diffraction so that the magnetic field could be applied along specific crystallographic directions. Both zero-field cooled (ZFC) and field-cooled (FC) protocols were used for the measurements. Samples were fixed to a Kel-F disk mounted inside a plastic straw; the magnetic signal of the mounting disk was measured separately and subtracted from the total moment.
[…]
@@ 3.1 Chemical and structural characterization
Table 1: Single crystal X-ray diffraction measurement and refinement information for Eu0.8Pt6Al16.4.
Chemical formula Eu0.781(5)Pt6Al16.4(1)
Formula Weight 1730.2
@@
Data/restraint/parameters 2204/1/100
Goodness-of-fit on $F^{2}$ 1.25
Final $R$ indexes $I\geq 2\sigma(I)$] $R_{1}=0.0207$, $wR_{2}=0.0208$
Final $R$ indexes [all data] $R_{1}=0.0387$, $wR_{2}=0.02245$
Largest diff. peak/hole 2.71 $e^{-}\text{\AA }^{-3}$/-2.26 $e^{-}\text{\AA }^{-3}$
Single-crystal X-ray diffraction measurement on the obtained crystals of Eu0.8Pt6Al16.4, revealed that it adopts a crystal structure Eu1-xPt6Al17-y ($x\sim 0.2$, $y\sim 0.6$) that does not match with any currently reported one. Table 1 shows the information on the measurement and refinement. The system is hexagonal and belongs to the space group $P\bar{6}2m$. A diagram of the unit cell basis of the structure is presented in Fig. 1, and the detailed information of the atomic positions, occupancies and thermal displacement parameters are given in Tab. 2. There are two distinct Eu sites ($2c$ and $2e$) located in the center of the blue and cyan polyhedra defined by the surrounding Pt and Al atoms in Fig. 1(a). The Eu $2e$ sites (in the cyan polyhedra) have an occupancy of around 56(1)%, whereas the other has full occupancy. Besides the Eu $2e$ position, also the Al $6i$ and Al $2e$ positions show an occupancy lower than 1, while all other positions are fully occupied within three standard deviations. The two Al $2e$ atoms form Al2 dumbbells that occupy the same space as the Eu $2e$ atoms, therefore only the Eu atom or the Al dumbbell can be present at the same time. This is also reflected by the occupancy factors of the Eu $2e$ (56%) and the Al dumbbell (46%), both obtained in a free refinement. The occupancy of the Al $2e$ positions was constrained to be equal. The resulting composition is therefore Eu0.781(5)Pt6Al16.4(1), corresponding to a $x=0.219(5)$ and a $y=0.6(1)$. Since the Eu occupation is close to 1/2, different symmetry reductions (subgroup $P3_{1}m$ of index 2 and subgroup $Cm2m$ of index 3 alongside the introduction of the repective twins/trillings) were tested, however, no conclusive structural model was obtained in a lower symmetric space group. When looking at the structure refinement and the site occupancies, an ordering might occur when changing the composition so that both the Eu $2e$ position and the Al dumbbell are occupied by 50%. This might lead to a doubling of the $c$-axis allowing an ordering of dumbbells and Eu atoms.
Table 2: Refined atomic positions, site occupancy, and isotropic thermal displacement parameters for Eu0.8Pt6Al16.4, according to SCXRD.
@@
Some of the crystals obtained were ground in order to measure powder X-ray diffraction, obtaining the pattern shown in Fig. 2, plotted with black symbols. A Rietveld refinement was performed, yielding $\chi^{2}=4.5$ and $R_{w}=15.5$. The model structure obtained from the single crystal X-ray refinement was used as the starting point; only the global factor, lattice parameters, atomic positions, microstrains, sample displacement and background were refined. The rest of the parameters, including the thermal displacements and the atomic fractions were kept fixed according to the values reported for the single crystal. The Rietveld-refined profile is plotted as a red curve, showing good agreement with the measured data, as also indicated by the difference (green line).
Figure 2: Powder X-ray diffraction of Eu0.8Pt6Al16.4 showing measured intensities (black points), the Rietveld-refined profile (red trace), and difference curve (green).
Table 3: Elemental analysis of Eu0.8Pt6Al16.4.
@@ 3.2 Electronic and magnetic characterization
Fig. 3 shows the ZFC and FC temperature-dependent magnetization normalized by the applied field of 50 Oe, applied along three mutually orthogonal directions: along the $a$-axis (in blue), along the $c$-axis (in green), and along the reciprocal lattice vector $b^{*}$ (in red). These three directions are schematically indicated with arrows on a prism drawing in the main panel. The inset shows an enlarged view of the low temperature range of these results, in order to see the feature more clearly in all three directions of applied field. The magnetization along $a$ and $b^{*}$ shows a downturn upon cooling below 3 K, whereas the magnetization along $c$ shows a more subtle kink and remains higher at low temperatures. The ZFC and FC data remain indistinguishable from each other throughout the measured temperature range.
Figure 3: Magnetization as a function of temperature for Eu0.8Pt6Al16.4 measured with 50 Oe along the $a$, $b^{*}$, and $c$ axes (blue, red and green, respectively). Data taken under zero-field- and field-cooled conditions overlap throughout the range. The inset expands the low-temperature region to emphasize the observed feature and includes the polycrystalline average (black).
@@
The magnetization, in units of Bohr magneton, $\mu_{B}$, per Eu atom is plotted as a function of the applied field along $a$, $b^{*}$ and $c$ in Fig. 4, measured at a constant temperature of 2 K with an increasing magnetic field. As was the case for Fig. 3, the formula unit of Eu0.781(5)Pt6Al16.4(1) was used for computing the intrinsic value of magnetization per Eu. The inset shows an enlarged view of the low-field data, showing a linear behavior that extrapolates to $M(H=0)=0$. No hysteresis is expected in any of the three directions, since no difference was observed between the temperature-dependent ZFC and FC measurements done at 50 Oe (see Fig. 3). As a reference, the black squares in the inset of Fig. 4 show the corresponding ZFC and FC magnetization values extracted from the $M(T)$ measurements at 2 K and a field of 50 Oe applied parallel to $c$, both consistent with the $M(H)$ results. The measurements along $a$ and $b^{*}$ both display some metamagnetic transition around 12 kOe, which is broadened due to the fact that the measurements were taken at a temperature that is close to the transition temperature. Combined, the $M(T)$ and $M(H)$ measurements shown in Figs. 3 and 4 indicate that the sample at low temperatures exhibits antiferromagnetic ordering, without signs of any ferromagnetic component.
@@
The magnetization shown in Fig. 4 does not reach full saturation at the maximum field measured, but it is consistent with a saturation of 7 $\mu_{B}$ expected for Eu2+ ions with $J=7/2$, indicated with a horizontal dashed line in the figure when using the proper stoichiometry of Eu0.781(5)Pt6Al16.4(1). The $\sim 5\%$ difference between the magnetization for $H||a$ (blue curve) and $H||b^{*}$ (red curve) could be due to slight displacements of the sample with respect to the central axis of the MPMS.
@@
Figure 4: Field-dependent magnetization at 2 K for Eu0.8Pt6Al16.4 with field applied along $a$ (blue), $b^{*}$ (red), and $c$ (green). The dashed horizontal line marks the moment expected for Eu2+ ($J=7/2$). Inset: low-field segment showing linear behavior and the field used for the temperature scans. The overlapping black squares show the corresponding ZFC and FC magnetization at 2 K obtained from the $M(T)$ measurements done with a field of 50 Oe applied parallel to $c$.
@@
Figure 5: Temperature evolution of the polycrystalline-averaged magnetic susceptibility $\chi_{\text{ave}}$ (black) and its derivative $d(\chi_{\text{ave}}T)/dT$ (magenta). The vertical dashed line identifies the Néel temperature $T_{N}=2.8(2)$ K obtained from the derivative maximum.
Figure 6: Resistivity normalized to the room-temperature value for Eu0.8Pt6Al16.4, measured with current along $a$ (blue) and $c$ (green) during cooling and warming cycles. Insets enlarge the low-temperature region; the black arrows indicate the estimated transition temperature according to $d(\chi_{\text{ave}}T)/dT$.
Figure 7: Magnetic susceptibility for Eu0.8Pt6Al16.4 measured with a 100 Oe field along $c$ (solid symbols, left axis) and its reciprocal (open symbols, right axis). The straight line represents a Curie–Weiss fit yielding $\mu_{eff}=8.3(3)\mu_{B}$ and $\theta_{CW}=–5.0(1)$ K, consistent with Eu2+ behavior.
@@
This yielded $\mu_{\text{eff}}=8.3(3)\ \mu_{B}$, which agrees with the expected value for Eu2+ ions (7.94 $\mu_{B}$) within uncertainty; and $\theta_{CW}=-5.0(1)$ K, which would indicate the paramagnetic phase has predominantly antiferromagnetic interactions, and is consistent with the low-temperature antiferromagnetic ordering.